OpenStudy (anonymous):

Kelly joins Powerhouse Gym. The cost is $30 per month and a one-time $100 joining fee. Lisa joins MegaMuscle Gym for $35 per month and a $50 joining fee. In how many months will Lisa have paid more than Kelly? A) 10 months B) 11 months C) 8 months D) 2 months

OpenStudy (anonymous):

set this up as a system of equations basically you have to realize that you want to create an equation for each girl for how much they pay. Kelly is K = 30*m + 100, where 'm' is the number of months Kelly has been part of the gym and 'K' is the amount Kelly has paid Lisa is L = 35*m + 50, where 'm' is the number of months Lisa has been part of the gym and 'L' is the amount Lisa has paid now set those two equations equations equal to each other to find out how many months it would take for Lisa to have paid as much as Kelly. so basically, solve: \[30m + 100 = 35m + 50\] for 'm'

OpenStudy (anonymous):

im really bad at this but is it 2 months

OpenStudy (anonymous):

its 11 months